OpenStudy (anonymous):

Question 7 (Multiple Choice Worth 5 points) [4_07]Choose the equation of the line passing through the point (-2, 6) and parallel to y = 1/2x - 8. y =1/2x + 7 y = 1/2x + 5 y = 2x + 7 y = 2x + 5

OpenStudy (anonymous):

@joemath314159

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

because it is parallel, the slope of your final equation is 1/2 plug in the point (-2,6) into y = (1/2)x + b to solve for b 6 = (1/2)(-2) + b 7 = b y = (1/2)x + 7, A

OpenStudy (experimentx):

any line parallel to y = 1/2x - 8 is y = 1/2x + c --------- 1 since it passes through, (-2, 6) 6 = 1/2(-2) + c find the value of c ... plugin in 1, and get you answer